1. A box of mass 5 kg is accelerated from rest across a floor at a rate of 2 m/s2  for 7 s. Find the net work done on the box.
1
Expert's answer
2021-12-27T08:01:07-0500

Given quantities:

"m = 5kg"

"a = 2 \frac{m}{s^2}"

"t = 7s"

"A= FS=mg\frac{at^2}{2}=5*10*\frac{2*49}{2}=2450J"